# Constraint on the physical origin of GRB prompt emission via its non-detected diffuse neutrino emission

### v1

@@ 4 Conclusion and discussion
The multi-messenger astronomy approach is a powerful tool to help us understand the physical origin of astronomical sources. GRBs are the most luminous explosions in the universe, and have rapidly advanced in recent years (Zhang, 2018). Especially, the jet composition of GRBs remains an open question (Zhang, 2011). The direct method to identify the jet composition of GRBs is to use the types of observed spectra for GRB prompt emission, but the uncertainty of measured is still very large (Zhang et al., 2011; Gao & Zhang, 2015; Li et al., 2024; Chang et al., 2024). In this paper, we propose an independent method to constrain the fraction of GRB prompt emission models based on the non-detected high-energy diffuse neutrino which is possibly associated with GRBs. For given the benchmark parameters of GRBs, i.e., $\Gamma=300$ and $\varepsilon_{p}\text{/}\varepsilon_{e}=10$, by adopting the method of summing up the individual GRB contributions, we find that the fractions of the dissipative photosphere model, the internal shock model, and the ICMART model are range in [0, $0.5\%$], [0, $1.1\%$], and $98.9\%$, 1], respectively. By adopting the method of assumed luminosity functions of GRB, the corresponding fractions of above three models are [0, $6.1\%$], [0, $8.2\%$], and $91.8\%$, 1], respectively. However, such fractions of different models are also dependent on the parameters of $\Gamma$ and $\varepsilon_{p}\text{/}\varepsilon_{e}$ (see Figure 2, 3, and Table 1, 2).
@@
Li et al. (2024) analyzed the time-resolved spectrum for all Fermi GRBs with known redshift to estimate the magnetization factor ($\sigma_{0}$), and found that the lower limits of $1+\sigma_{0}$ for $22.7\%$ of the non-thermal spectrum (225 out of 318) are larger than 10. By adopting the benchmark parameters of $\Gamma=300$ and $\varepsilon_{p}\text{/}\varepsilon_{e}=10$ in our calculations, we estimate the lower limit of the fraction of the ICMART model and conclude that most of GRBs should originate from the ICMART model. Our results are not contradictory to those from Li et al. (2024) who obtained the lower limit of the magnetization factor. This suggests that most GRB jets should contain a significant magnetic-energy component, namely, magnetic dissipation. On the other hand, Aartsen et al. (2017) suggested that the internal shock model and the photospheric model are excluded at the $99\%$ confidence level for benchmark parameters of $\Gamma$ and $\varepsilon_{p}\text{/}\varepsilon_{e}$ when they assumed that the prompt emission of all observed GRBs originates from a single mode. However, by taking into account a certain fraction of the GRB model that is from the ICMART model, it is found that both the internal shock model and the photospheric model can survive for the benchmark parameters.
@@
Our analysis also poses a curious question. The fact that the lack of most GRBs redshift measured by adopting the method of summing up the individual GRB contributions, we have to adopt an average values of redshift distribution for those GRBs which is also adopted by the IceCube collaboration (Aartsen et al., 2017). In order to test the effect of varying redshift, we also adopt different values of redshift, and the predicted flux of diffuse neutrino for the benchmark parameters base on different redshift value assumption are shown in Figure 4. In the left panel of Figure 4, we test the effect of changing mean redshift of short GRBs $z=0.3$, 0.5, and 1.0 for fixed $z=2.15$ of long GRBs. The middle panel shows the effect of changing the mean redshift of long GRBs $z=1.0$, 2.0, and 3.0 for fixed $z=0.5$ of short GRBs, and the right panel in Figure 4 shows the result of simultaneously increasing the redshifts of both long and short GRBs. We find that changing the mean redshift of short GRBs has a little impact on the results. That is because the fraction of short GRBs in our sample is not larger enough, and the isotropic energy of short GRBs is usually lower than that of long GRBs. It means that the ability to produce neutrinos is also weaker by comparing the long GRBs. Moreover, it is found that the flux of diffuse neutrino is increasing with the increasing redshift. The reason is possible that increasing redshift can also result in increasing $L_{\rm iso}$ for a given $F_{\gamma}$. In any case, increased $L_{\rm iso}$ could lead to increase of $p\gamma$ optical depth, and increase the pion production efficiency if pion production efficiency not saturated. We also find that the effect of diffuse neutrino flux with varying redshifts of GRBs is significant larger for the large dissipation radius model (i.e. ICMART model). That is because the pion production efficiency is already saturated in most GRBs in smaller dissipation radius model (i.e. dissipative photosphere model), and further increases in $p\gamma$ optical depth do not increase pion production efficiency. In the future, with the increasing number of GRB redshift measurements, the constraint on the fraction of GRB prompt emission models from the effect of redshift could be ignored.
@@
On the other hand, it is worth noting that the constraints of this work strongly depend on the dissipation radius of photosphere, internal shock, and the ICMART models we adopted. In this work, we only adopt the typical values of dissipation radius to constrain the physical origin of GRB prompt emission (Zhang, 2018), and more precise constraints will be obtained once the dissipation radius can be accurately determined.

### v2

@@ 4 Conclusion and discussion
The multimessenger astronomy approach is a powerful tool to help us understand the physical origin of astronomical sources. GRBs are the most luminous explosions in the Universe and have rapidly advanced in recent years (2018pgrb.book.....Z). Especially, the jet composition of GRBs remains an open question (2011CRPhy..12..206Z). The direct method to identify the jet composition of GRBs is to use the types of observed spectra for GRB prompt emission, but the uncertainty of the measurements is still very large (2011ApJ...730..141Z; 2015ApJ...801..103G; 2024ApJ...972....1L; 2024ApJS..275....9C). In this paper, we propose an independent method to constrain the fraction of GRB prompt emission models based on the nondetected high-energy diffuse neutrino, which is possibly associated with GRBs. For given the benchmark parameters of GRBs, i.e., $\Gamma=300$ and $\varepsilon_{p}\text{/}\varepsilon_{e}=10$, by adopting the method of summing up the individual GRB contributions, we find that the fractions of the dissipative photosphere model, the internal shock model, and the ICMART model are range in [0, $0.5\%$], [0, $1.1\%$], and $98.9\%$, 1], respectively. By adopting the method of assumed luminosity functions of GRBs, the corresponding fractions of above three models are [0, $6.1\%$], [0, $8.2\%$], and $91.8\%$, 1], respectively. However, such fractions of different models are also dependent on the parameters of $\Gamma$ and $\varepsilon_{p}\text{/}\varepsilon_{e}$ (see Figure 2 and 3 and Table 1 and 2).
@@
2024ApJ...972....1L analyzed the time-resolved spectrum for all Fermi GRBs with known redshift to estimate the magnetization factor ($\sigma_{0}$) and found that the lower limits of $1+\sigma_{0}$ for $22.7\%$ of the nonthermal spectrum (225 out of 318) are larger than 10. By adopting the benchmark parameters of $\Gamma=300$ and $\varepsilon_{p}\text{/}\varepsilon_{e}=10$ in our calculations, we estimate the lower limit of the fraction of the ICMART model and conclude that most of GRBs should originate from the ICMART model. Our results are not contradictory to those from 2024ApJ...972....1L, who obtained the lower limit of the magnetization factor. This suggests that most GRB jets should contain a significant magnetic-energy component, namely, magnetic dissipation. On the other hand, 2017ApJ...843..112A suggested that the internal shock model and the photospheric model are excluded at the $99\%$ confidence level for benchmark parameters of $\Gamma$ and $\varepsilon_{p}\text{/}\varepsilon_{e}$ when they assumed that the prompt emission of all observed GRBs originates from a single mode. However, by taking into account a certain fraction of the GRB model that is from the ICMART model, it is found that both the internal shock model and the photospheric model can survive for the benchmark parameters.
@@
Our analysis also poses a curious question. To calculate the diffuse neutrino flux by summing up individual GRB contributions, we have to adopt average redshift values for the majority of GRBs that lack redshift measurements. This approach of using average values is also employed by the IceCube collaboration (2017ApJ...843..112A). In order to test the effect of varying redshift, we also adopt different values of redshift, and the predicted flux of diffuse neutrino for the benchmark parameters based on different redshift value assumption is shown in Figure 4. In the left panel of Figure 4, we test the effect of changing mean redshift of short GRBs $z=0.3$, 0.5, and 1.0 for fixed $z=2.15$ of long GRBs. The middle panel shows the effect of changing the mean redshift of long GRBs $z=1.0$, 2.0, and 3.0 for fixed $z=0.5$ of short GRBs, and the right panel in Figure 4 shows the result of simultaneously increasing the redshifts of both long and short GRBs. We find that changing the mean redshift of short GRBs has a little impact on the results. That is because the fraction of short GRBs in our sample is not large enough, and the isotropic energy of short GRBs is usually lower than that of long GRBs. It means that the ability to produce neutrinos is also weaker by comparing the long GRBs. Moreover, it is found that the flux of diffuse neutrino is increasing with the increasing redshift. The reason is possibly that increasing redshift can also result in increasing $L_{\rm iso}$ for a given $F_{\gamma}$. In any case, increased $L_{\rm iso}$ could lead to an increase of $p\gamma$ optical depth and an increase of the pion production efficiency if it is not saturated. We also find that the effect of diffuse neutrino flux with varying redshifts of GRBs is significantly larger for the large dissipation radius model (i.e. ICMART model). That is because the pion production efficiency is already saturated in most GRBs in a smaller dissipation radius model (i.e., the dissipative photosphere model), and further increases in $p\gamma$ optical depth do not increase pion production efficiency. In the future, with the increasing number of GRB redshift measurements, the constraint on the fraction of GRB prompt emission models from the effect of redshift could be ignored.
@@
On the other hand, it is worth noting that the constraints of this work strongly depend on the dissipation radius of photosphere, internal shock, and the ICMART models we adopted. In this work, we only adopt the typical values of the dissipation radius to constrain the physical origin of the GRB prompt emission (2018pgrb.book.....Z), and more precise constraints will be obtained once the dissipation radius can be accurately determined.
@@ Acknowledgments
Figure 1: Predicted flux of diffuse neutrino for the benchmark parameters of $\varepsilon_{p}\text{/}\varepsilon_{e}=10$ and $\Gamma=300$ by assuming that all GRBs are originated from the signal model (e.g., the dissipative photosphere model, the internal shock model, and the ICMART model). The bands for the diffuse neutrino flux are displayed for $\varepsilon_{p}\text{/}\varepsilon_{e}=3-10$ and $\Gamma=100-500$. The gray line with arrows is the upper limit flux of diffuse neutrino. Left: results using the method of summing up the individual GRB contributions. Right: results using the method of assumed luminosity functions of the GRB.
Figure 2: Constraint on the fraction of GRB prompt emission models by adopting the method of summing up the individual GRB contributions. Left: the constraint results for the varying value of $\Gamma=$ 300 and 500 with fixed $\varepsilon_{p}\text{/}\varepsilon_{e}=10$. Right: the constraint results for the varying value of $\varepsilon_{p}\text{/}\varepsilon_{e}=3$, 5, and 10 with fixed $\Gamma=300$.
Figure 3: Constraint on the fraction of GRB prompt emission models by adopting the method of assumed luminosity functions of GRB. Left: the constraint results for the varying value of $\Gamma=$ 300 and 500 with fixed $\varepsilon_{p}\text{/}\varepsilon_{e}=10$. Right: the constraint results for the varying value of $\varepsilon_{p}\text{/}\varepsilon_{e}=$ 5 and 10 with fixed $\Gamma=300$.
Figure 4: Predicted flux of diffuse neutrino for the benchmark parameters of $\varepsilon_{p}\text{/}\varepsilon_{e}=10$ and $\Gamma=300$ by assuming that all GRBs are originated from signal model with different mean redshift. Left: changed the mean redshift of short GRBs as $z=0.3$, 0.5, and 1.0 for fixed $z=2.15$ of long GRBs. Middle: changed the mean redshift of long GRBs as $z=1.0$, 2.0, and 3.0 for fixed $z=0.5$ of short GRBs. Right: simultaneously increasing the redshift of long GRBs ($z=1.0$, 2.0, and 3.0) and short GRBs ($z=0.25$, 0.5, and 1.0).
Table 1: The fraction range of the photosphere model, the internal shock model, and the ICMART model within different $\varepsilon_{p}\text{/}\varepsilon_{e}$ and $\Gamma$ by adopting the method of summing up the individual GRB contributions.
$\varepsilon_{p}\text{/}\varepsilon_{e}$ $\Gamma$ Range of Fraction
Photosphere Internal Shock ICMART
10 100 Rule out Rule out Rule out
300 [0, 0.5%] [0, 1.1%] [98.9%, 1]
500 [0, 7.1%] [0, 31.7%] [68.3%, 1]
3 300 [0, 44.9%] No constraint No constraint
5 [0, 19.3%] [0, 53.1%] [46.9%, 1]
10 [0, 0.5%] [0, 1.1%] [98.9%, 1]
Table 2: The fraction range of the photosphere model, the internal shock model, and the ICMART model within different $\varepsilon_{p}\text{/}\varepsilon_{e}$ and $\Gamma$ by adopting the method of assumed luminosity functions of GRB.
$\varepsilon_{p}\text{/}\varepsilon_{e}$ $\Gamma$ Range of Fraction
Photosphere Internal Shock ICMART
10 100 Rule out Rule out Rule out
300 [0, 6.1%] [0, 8.2%] [91.8%, 1]
500 [0, 25.5%] [0, 48.0%] [52.0%, 1]
3 300 No constraint No constraint No constraint
5 [0, 78.5%] No constraint No constraint
10 [0, 6.1%] [0, 8.2%] [91.8%, 1]